# A Superconducting Single-Atom Phonon Laser

### v2

@@ CONCLUSION
This article demonstrates the experimental realization of a phononic analog to the optical single-atom laser. Our experiment consists of a superconducting single-atom, realized using a transmon style qubit coupled resonantly to an HBAR mode. When driving on resonance with the HBAR mode, the intrinsic non-linearity of the qubit-phonon coupling generates a highly excited phonon state — the phonon laser. A key feature of this experiment is the driven two-level atom rather than the parametric instability driving phonon excitations. Moreover, the phonon mode is confined in a bulk longitudinal mode within a sapphire substrate. The bulk nature may allow the phonon laser’s integration with additional on-chip architectures, such as color centers or quantum dots.
@@
Our results have demonstrated the successful generation of a coherent phonon laser, and we have further demonstrated a counterintuitive feature of phonon lasers, an upper instead of a lower lasing threshold Ashhab et al. (2009). The size of the coherent lasing state achieved in this work was limited by both the qubit and phonon linewidth. However, the primary limiting factor was the phonon linewidth. Decreasing the decay rates will reduce the upper lasing threshold power and increase the peak lasing amplitude. Moreover, it should be possible to construct a single-atom microwave photon laser using a superconducting qubit weakly coupled to a microwave resonator in conventional cQED. Future studies could include a linear probe or use a higher-order transition of the transmon to drive and measure the phonon statistics. Such a cQAD-compatible phonon laser promises to provide a highly coherent source of phonons, which have applications including sensing to quantum information processing and the generation of high-displacement, Schrödinger cat states.
ACKNOWLEDGMENTS
The authors thank Enrique Sahagun for the device rendering Sci . W.J.M.F and G.A.S. acknowledge support through the QUAKE project, project number 680.92.18.04, of the research programme Natuurkunde Vrije Programma’s of the Dutch Research Council (NWO). C.A.P. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). A.M and V.A.S.V.B. acknowledge financial support from the Contrat Triennal 2021-2023 Strasbourg Capitale Europeenne.
Authors contributions C.A.P. performed experiments, theoretical modelling, conceptualization, and wrote the manuscript with input from all authors. W.J.M.F. fabricated the device and performed experiments. V.A.S.V.B. performed theoretical modelling. A.M. provided supervision and funding acquisition. G.A.S. provided supervision, conceptualization and funding acquisition.
Ethics declarations:
Competing interests: The authors declare no competing interests.
Data and materials availability All data, analysis code, and measurement software are available in the manuscript or the supplementary material or are available at Zenodo .
@@ Methods
For the gated two-tone measurement, port one of the VNA was set up just as in the two-tone measurement. An external signal generator generated the qubit drive tone. The qubit drive tone was passed through an RF switch before being combined with the readout tone using a directional coupler. The RF switch was triggered using a pulse generator at a 3 kHz repetition rate. The signal generator was set to a 22.5% duty cycle, so the qubit drive was off for 75 $\mu$s per trace. The VNA was synchronously triggered by the signal generator, allowing for 65536 trace averages to be performed (maximum setting). Each trace averaged measurement was repeated 75 additional times to improve the signal-to-noise ratio further.
Supplementary Information
I Device
Figure S1: Fabricated flip-chip HBAR Device. (A) Optical micrograph of the assembled flip-chip device. The top feedline of qubits is fabricated with a flip-chip of sapphire, and the bottom feedline has no sapphire flip-chip as references. (B) Zoomed optical micrograph of the flip-chip assembly. (C) Optical micrograph of the qubit with the sapphire chip assembled on top. The overlap between the antenna and the $250\mu$m aluminum nitride piezoelectric transducer is visible. The scale bar is $250\mu$m. (D) Optical micrograph of the entire chip loaded in the printed circuit board.
Figure S2: Schematic of the measurement setup. To the left is the general Dilution refrigerator wiring setup. Outside the refrigerator, we used two different setups. In (b), we show our Two-tone spectroscopy configuration used for qubit spectroscopy. Here, we sent a weak continuous wave tone (readout resonator probe) from the vector network analyzer (VNA) port one and a second continuous wave tone (qubit drive) from the VNA port 3. These two signals are combined using a directional coupler before entering the dilution refrigerator. The signal from the dilution refrigerator goes through a room-temperature amplifier before it goes to port 2 of the VNA. In (c), we show our ”time domain” setup. Here, we replaced output port 3 of the VNA with a signal generator (SG) to provide the drive tone. A switch (Taylor MW switch) together with a pulse generator (PG, Rigol DG1022) is placed between the directional coupler and the signal generator.
II Theory
@@ II.2 Mean field theory for the steady-state phonon population
The equation for the phonon number (S12) is a non-linear equation for $\bar{b}$ that has to be solved numerically. The results for parameters in correspondence with the experiment are shown in Fig.  S3, in which we can see a good agreement with the full master equation simulation.
Figure S3: Steady-state mean-field phonon number as a function of the qubit drive power. The red line depicts the result obtained with the mean-field solution of Eq. (S14), and the gray points depict the results from simulations of the full master equation (S5). The parameters used are given in Table 1 in correspondence with the experimental parameters of the main text.
@@
From such a mean-field perspective, the steady-state of the phonon mode behaves similarly as if it would be coupled to a coherently driven cavity with the nonlinear decays and detunings given by Eq.  (S15). We notice that as the power is increased, the effective decay $\gamma_{\rm{eff}}$ increases, a consequence of the qubit power broadening. Such an effect yields the eventual decoupling between the qubit and the phonon, similar to an inverse Purcell effect. Nevertheless, such a power broadening is not the only physical mechanism that plays a role in the build-up of the phonon steady-state. In fact, both the effective decay and detuning depend on the phonon amplitude $\bar{b}$, a non-linearity that can not be discarded for the experimental parameters. To show this effect, we plot in Fig. S4 the mean-field solution for $n_{\rm{b}}$ (red curve) and the corresponding curve by setting $g_{\rm{qb}}=0$ in Eqs. (S15) (dashed black curve). The latter still captures the inverse Purcell effect due to the qubit power broadening but yields half of the maximum phonon occupancy, indicating the importance of the phonon nonlinearity stemming from the phonon-qubit coupling and the intrinsic two-level nature of the qubit.
Figure S4: Steady-state mean-field phonon number as a function of the qubit drive power. The red line depicts the result obtained with the mean-field solution of Eq. (S14), while the black dashed line depicts the result obtaining by discarding the phonon nonlinearity in eqs. (S15) The parameters used are given in Table 1 in correspondence with the experimental parameters of the main text.
II.3 Qubit Two-Tone Spectroscopy
@@
In the above equations, $\omega^{(j)}_{\rm q}$ and $\Gamma^{(j)}_{\rm q}$ are the frequency and linewidth of the qubit with the readout resonator in the state $|j\rangle$, respectively. The intrinsic qubit linewidth, with the readout resonator in its ground state, is given by $\gamma_{\rm q}=\Gamma_{1}/2+\Gamma_{\phi}$, where $\Gamma_{1}$ is the longitudinal relaxation rate, and $\Gamma_{\phi}$ is the pure dephasing rate. The qubit frequency $\omega^{(0)}_{\rm q}$ is ac Stark shifted by $B$ from its intrinsic value $\omega_{\rm q}$. We have also assumed that the readout drive is on resonance with the readout resonator, i.e. $\Delta_{\rm r}=0$, and the readout resonator has a full-width half-maximum linewidth $\kappa$. In the limit $\chi\sim\kappa$ and with $\Delta_{\rm r}=0$, the components $S_{j}(\omega)$ have non-Lorentzian lineshapes and can even be negative. The sum of these individual components can result in an asymmetry of the qubit spectrum, as seen in Fig. S5(a).
@@
Figure S5: Two-tone qubit spectroscopy measurements. (A) Experimental two-tone qubit spectrum (blue) and the qubit spectrum determined from a fit to Eq. S19 (red). The qubit drive power at room temperature was set to -20.0 dBm. (B) Extracted value of the readout resonator photon population as a function of qubit drive power.
An example spectrum and its fit are shown in Fig. S5(a). First, the value of the dispersive shift was determined to be $\chi=-1.2\pm 0.2$ MHz, which agreed with our designed value. This value was then fixed, and the data was fit for all qubit drive powers to determine the remaining values. We found that the ground state readout photon population depended on the qubit drive power. This is likely due to the heating of the silicon substrate because the qubit drive tone was applied via the readout resonator. The value of $\bar{n}_{\rm g}(\mathcal{P}_{\rm d})$ is shown in Fig. S5(b). The zero power qubit linewidth was also extracted by extrapolating the measured power-broadened qubit linewidth to zero power and was determined to be $\gamma_{\rm q}(0)=0.420\pm 0.04$ MHz, which agrees with our $\gamma_{\rm q}=\Gamma_{1}/2+\Gamma_{\phi}\geq 1/2T_{1}$ limit determined using a time-domain measurement with $T_{1}=180\pm 30$ ns. The comparison between our spectroscopy and time domain measurements indicates that our qubit decay is dominated by decoherence, and therefore we have ignored intrinsic dephasing in our model since qubit dephasing will be dominated by power-induced dephasing induced by the qubit drive tone.
II.4 Master Equation Simulations
@@
Figure S6: Two-tone qubit spectroscopy measurements. (A) Experimental two-tone qubit spectrum (blue) overlapped with the qubit spectrum determined from the master equation simulation (red). The frequency is defined as the detuning from the bare qubit frequency. The qubit drive power at room temperature was set to -12.0 dBm. (B) Qubit spectrum and master equation simulation for a qubit drive power of -6.0 dBm. (C) Qubit spectrum and master equation simulation for a qubit drive power of -3.0 dBm. The feature at $\sim 0$ MHz is an additional HBAR mode, one free-spectral range from the HBAR mode of interest for this work. The only parameter that was varied within the simulation was the room-temperature value of the qubit drive.
@@
where $\delta$ calibrates the room-temperature power to the corresponding value of $\mathcal{P}_{\rm d}$ accounting for all losses and multiplicative factors. The value of $\delta$ was determined by matching the simulation to the power-broadened qubit spectrum at multiple qubit drive powers. A set of qubit spectra is shown in Fig. S6 where it can be seen that the master equation simulation is in excellent agreement capturing the qubit asymmetry at low power, and the power-broadened qubit spectrum at higher drive powers. It should be noted the measured transmission signal $S_{21}$ is proportional to the qubit population $\langle\hat{\sigma}_{\rm z}\rangle$ only if the appropriate signal quadrature is measured. Here, this was determined by calculating the rotation angle that minimized the signal in the out-of-phase quadrature. To confirm, we compared our single-trace two-tone data with a direct measurement of the readout resonator frequency at several qubit drive powers. At the highest drive powers, we observed a slight deviation resulting from signal mixing into the out-of-phase quadrature. However, this was at the drive powers above the self-quenching threshold and resulted in a slight mismatch between our experimental and simulated two-tone traces.
@@
Figure S7: Gated two-tone qubit spectroscopy. (A) Experimental gated two-tone measurements for qubit drive powers -12.0, -6.0, -3.0 dBm. The master equation simulations at the same drive power are plotted as dashed black lines. The simulation has an excellent agreement for the ringdown data for all powers and a slight deviation from the ring-up data. This could be attributed to higher-order non-linearities we are not considering within our simple master equation model. However, the agreement between the experiment and simulation is good for all powers. (B) Simulated phonon state population $\langle\hat{b}^{\dagger}\hat{b}\rangle$ as a function of qubit drive power. (C) Simulated second-order phonon correlations $g^{(2)}(0)$ as a function of qubit drive power.
@@
Following the calibration of the qubit spectrum, the phonon mode was included in the master equation simulation. The value of the coupling rate $g_{\rm qb}$ and the phonon linewidth $\gamma_{\rm b}$ were determined by performing a fit to the time domain gated two-tone measurements. The fit was determined by fitting both the ringdown (as shown in the main text) and also the ring-up of the qubit. Here, the device was allowed to thermalize before the gated-ringdown measurement, weakly probing the readout resonator continually. At the time $t=0$, the qubit drive is switched on for $175\mu$s, allowing the phonon moded to ring-up. At $t=175\mu$s, the qubit drive is switched off, and the phonon ringdown is observed, as described in the main text. The best-fit values were compared at multiple drive powers; see Fig. S7(a). We extracted a value of the qubit phonon coupling of $g_{\rm qb}/2\pi=162$ kHz and a phonon linewidth of $\gamma_{\rm b}/2\pi=6.81$ kHz. From this set of simulations, the steady-state phonon population and phonon statistics can be estimated for multiple drive powers. We observe a good agreement between the numerical simulation and the ring-up data and an excellent agreement between the ringdown data for all powers. Deviations in the ring-up simulations likely result from higher-order nonlinearities we are not considering in our model.
@@
The master equation simulations directly calculate the time dynamics of $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$, this value has been scaled by the same constant factor for all powers to give a direct comparison to our two-tone measurement. Since we expect the variation in the transmission coefficient measured in the two-tone measurement to be proportional to $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$, see Eq. S3 and Ref. Gambetta et al. (2006). From our simulation, we calculate the phonon population and second-order correlation function for multiple drive powers, shown in Fig. S7(b,c). For low drive powers, the phonon statistics are described by a coherent state, $g^{(2)}(0)\approx 1.0$, as expected for a laser. However, as discussed in the main text, single-atom lasers exhibit self-quenching above a given upper-threshold power. This can be seen in both the phonon population, as a rapid decrease in the populations, and in the phonon statistics as $g^{(2)}(0)>1.0$ for drive powers above the self-quenching threshold.
Table 1: Symbols and parameters
@@
III Simplified Single-Atom Laser Model

### v3

@@ Discussion
This article demonstrates the experimental realization and generation of large-amplitude phonon states in cQAD. Our experiment consists of a superconducting single-atom, realized using a transmon-style qubit coupled resonantly to an HBAR mode. When driving on resonance with the HBAR mode, the intrinsic non-linearity of the qubit-phonon coupling generates a highly excited phonon state. A key feature of this experiment is the driven two-level atom rather than the parametric instability driving phonon excitations. Moreover, the phonon mode is confined in a bulk longitudinal mode within a sapphire substrate. The bulk nature may allow the phonon mode’s integration with additional on-chip architectures, such as color centers or quantum dots.
@@
Our results have demonstrated the successful generation of a large amplitude phonon state, and we have further demonstrated a counterintuitive feature predicted by our model, an upper threshold Ashhab et al. (2009). The size of the coherent state achieved in this work was limited by both the qubit and phonon linewidth. However, the primary limiting factor was the phonon linewidth. Decreasing the decay rates will reduce the upper threshold power and increase the peak phonon amplitude. Future studies could include a linear probe or use a higher-order transition of the transmon to drive and measure the phonon statistics. Such a cQAD-compatible system promises to provide a highly coherent source of phonons, which have applications including sensing to quantum information processing and the generation of high-displacement, Schrödinger cat states.
@@ Methods
For the gated two-tone measurement, port one of the VNA was set up just as in the two-tone measurement. An external signal generator generated the qubit drive tone. The qubit drive tone was passed through an RF switch before being combined with the readout tone using a directional coupler. The RF switch was triggered using a pulse generator at a 3 kHz repetition rate. The signal generator was set to a 22.5% duty cycle, so the qubit drive was off for 75 $\mu$s per trace. The VNA was synchronously triggered by the signal generator, allowing for 65536 trace averages to be performed (maximum setting). Each trace-averaged measurement was repeated 75 additional times to improve the signal-to-noise ratio further.
Data Availability
All data, analysis code, and measurement software are available in the manuscript or the supplementary material or are available at Zenodo https://doi.org/10.5281/zenodo.14810526.
Acknowledgments
The authors thank Enrique Sahagun for the device rendering Sci . W.J.M.F and G.A.S. acknowledge support through the QUAKE project, project number 680.92.18.04, of the research programme Natuurkunde Vrije Programma’s of the Dutch Research Council (NWO). C.A.P. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) (PDF-567689-2022). A.M and V.A.S.V.B. acknowledge financial support from the Contrat Triennal 2021-2023 Strasbourg Capitale Europeenne.
Author contributions C.A.P. performed experiments, theoretical modelling, conceptualization, and wrote the manuscript with input from all authors. W.J.M.F. fabricated the device and performed experiments. V.A.S.V.B. performed theoretical modelling. A.M. provided supervision and funding acquisition. G.A.S. provided supervision, conceptualization and funding acquisition.
Ethics declarations:
Competing interests: The authors declare no competing interests.
Supplementary Information: Large Amplitude Phonon States in Quantum Acoustics
[…]
Email: clinton.potts@nbi.ku.dk
Email: g.a.steele@tudelft.nl
I Supplementary Note 1: The Device
SUPPLEMENTARY FIGURE 1: Fabricated flip-chip HBAR Device. (A) Optical micrograph of the assembled flip-chip device. The top feedline of qubits is fabricated with a flip-chip of sapphire, and the bottom feedline has no sapphire flip-chip as references. (B) Zoomed optical micrograph of the flip-chip assembly. (C) Optical micrograph of the qubit with the sapphire chip assembled on top. The overlap between the antenna and the $250\mu$m aluminum nitride piezoelectric transducer is visible. The scale bar is $250\mu$m. (D) Optical micrograph of the entire chip loaded in the printed circuit board.
SUPPLEMENTARY FIGURE 2: Schematic of the measurement setup. (A) Dilution refrigerator wiring setup. Outside the refrigerator, we used two different setups. In (B), we show our Two-tone spectroscopy configuration used for qubit spectroscopy. Here, we sent a weak continuous wave tone (readout resonator probe) from the vector network analyzer (VNA) port one and a second continuous wave tone (qubit drive) from the VNA port 3. These two signals are combined using a directional coupler before entering the dilution refrigerator. The signal from the dilution refrigerator goes through a room-temperature amplifier before it goes to port 2 of the VNA. In (C), we show our ”time domain” setup. Here, we replaced output port 3 of the VNA with a signal generator (SG) to provide the drive tone. A switch (Taylor MW switch) together with a pulse generator (PG, Rigol DG1022) is placed between the directional coupler and the signal generator. Adapted from Franse et al. (2024)
II Supplementary Note 2: Theory
@@ II.2 Mean field theory for the steady-state phonon population
The equation for the phonon number (S12) is a non-linear equation for $\bar{b}$ that has to be solved numerically. The results for parameters in correspondence with the experiment are shown in Supplementary Fig.  3, in which we can see a good agreement with the full master equation simulation.
SUPPLEMENTARY FIGURE 3: Steady-state mean-field phonon number as a function of the qubit drive power. The red line depicts the result obtained with the mean-field solution of Eq. (S14), and the gray points depict the results from simulations of the full master equation (S5). The parameters used are given in Table 1 in correspondence with the experimental parameters of the main text.
@@
From such a mean-field perspective, the steady-state of the phonon mode behaves similarly as if it would be coupled to a coherently driven cavity with the nonlinear decays and detunings given by Eq.  (S15). We notice that as the power is increased, the effective decay $\gamma_{\rm{eff}}$ increases, a consequence of the qubit power broadening. Such an effect yields the eventual decoupling between the qubit and the phonon, similar to an inverse Purcell effect. Nevertheless, such a power broadening is not the only physical mechanism that plays a role in the build-up of the phonon steady-state. In fact, both the effective decay and detuning depend on the phonon amplitude $\bar{b}$, a non-linearity that can not be discarded for the experimental parameters. To show this effect, we plot in Fig. 4 the mean-field solution for $n_{\rm{b}}$ (red curve) and the corresponding curve by setting $g_{\rm{qb}}=0$ in Eqs. (S15) (dashed black curve). The latter still captures the inverse Purcell effect due to the qubit power broadening but yields half of the maximum phonon occupancy, indicating the importance of the phonon nonlinearity stemming from the phonon-qubit coupling and the intrinsic two-level nature of the qubit.
SUPPLEMENTARY FIGURE 4: Steady-state mean-field phonon number as a function of the qubit drive power. The red line depicts the result obtained with the mean-field solution of Eq. (S14), while the black dashed line depicts the result obtaining by discarding the phonon nonlinearity in eqs. (S15) The parameters used are given in Table 1 in correspondence with the experimental parameters of the main text.
II.3 Phonon mode linewidth
The semi-classical equations describing the system dynamics (S11) can also be used to obtain information about the phonon mode response in the steady-state, in particular, its effective linewidth. To compute the effective phonon linewidth in the steady-state, we numerically solve Eq. (S11) in time-domain, which gives the time-dependent phonon amplitude $b(t)$. The phonon spectrum is then given by $|b[\omega]|^{2}$, where $b[\omega]$ is the Fourier transform of $b(t)$. Such a spectrum is a Lorentzian in the frequency domain, with a linewidth that we refer as the Phonon linewidth.
In Supplementary Fig. 5 we shown in (A) the steady-state phonon population as a function of the drive amplitude for the parameters of the experiment and in (B) the corresponding phonon linewdith obtained by fitting the phonon spectrum obtained numerically to a Lorentzian. In (C) and (D) we show two examples of phonon spectrum. We notice that as the drive approaches the upper threshold, the phonon linewidth goes below its intrinsic value, pointing to an amplification of the phonon mode. Such an effect would not be possible only with a decoupling between qubit and phonon, and it is intrinsically related to the non-linear character of the Jaynes-Cummings interaction. Furthermore, the reduction of the effect linewidth is what enables the build-up of such a large steady-state phonon population.
SUPPLEMENTARY FIGURE 5: Linewidth narrowing of the HBAR mode (A) Steady-state phonon occupation for the system in the manuscript. (B) Effective phonon linewidth extracted from a Lorentzian fit of the phonon spectrum. (C,D) phonon spectra at two representative drive amplitudes showing the linewidth narrowing with power, the black dashed line depicts the spectrum for a phonon mode uncoupled from the two-level system. Parameters in correspondence with the experiment.
II.4 Qubit Two-Tone Spectroscopy
@@
In the above equations, $\omega^{(j)}_{\rm q}$ and $\Gamma^{(j)}_{\rm q}$ are the frequency and linewidth of the qubit with the readout resonator in the state $|j\rangle$, respectively. The intrinsic qubit linewidth, with the readout resonator in its ground state, is given by $\gamma_{\rm q}=\Gamma_{1}/2+\Gamma_{\phi}$, where $\Gamma_{1}$ is the longitudinal relaxation rate, and $\Gamma_{\phi}$ is the pure dephasing rate. The qubit frequency $\omega^{(0)}_{\rm q}$ is ac Stark shifted by $B$ from its intrinsic value $\omega_{\rm q}$. We have also assumed that the readout drive is on resonance with the readout resonator, i.e. $\Delta_{\rm r}=0$, and the readout resonator has a full-width half-maximum linewidth $\kappa$. In the limit $\chi\sim\kappa$ and with $\Delta_{\rm r}=0$, the components $S_{j}(\omega)$ have non-Lorentzian lineshapes and can even be negative. The sum of these individual components can result in an asymmetry of the qubit spectrum, as seen in Supplementary Fig. 6(a).
@@
SUPPLEMENTARY FIGURE 6: Two-tone qubit spectroscopy measurements. (A) Experimental two-tone qubit spectrum (blue) and the qubit spectrum determined from a fit to Eq. S19 (red). The qubit drive power at room temperature was set to -20.0 dBm. (B) Extracted value of the readout resonator photon population as a function of qubit drive power.
An example spectrum and its fit are shown in Supplementary Fig. 6(a). First, the value of the dispersive shift was determined to be $\chi/2\pi=-1.2\pm 0.2$ MHz, which agreed with our designed value. This value was then fixed, and the data was fit for all qubit drive powers to determine the remaining values. We found that the ground state readout photon population depended on the qubit drive power. This is likely due to the heating of the silicon substrate because the qubit drive tone was applied via the readout resonator. The value of $\bar{n}_{\rm g}(\mathcal{P}_{\rm d})$ is shown in Supplementary Fig. 6(b). The zero power qubit linewidth was also extracted by extrapolating the measured power-broadened qubit linewidth to zero power and was determined to be $\gamma_{\rm q}(0)/2\pi=0.420\pm 0.04$ MHz, which agrees with our $\gamma_{\rm q}/2\pi=\Gamma_{1}/2+\Gamma_{\phi}\geq 1/2T_{1}$ limit determined using a time-domain measurement with $T_{1}=180\pm 30$ ns. The comparison between our spectroscopy and time domain measurements indicates that our qubit decay is dominated by decoherence, and therefore we have ignored intrinsic dephasing in our model since qubit dephasing will be dominated by power-induced dephasing induced by the qubit drive tone.
II.5 Master Equation Simulations
@@
SUPPLEMENTARY FIGURE 7: Two-tone qubit spectroscopy measurements. (A) Experimental two-tone qubit spectrum (blue) overlapped with the qubit spectrum determined from the master equation simulation (red). The frequency is defined as the detuning from the bare qubit frequency. The qubit drive power at room temperature was set to -12.0 dBm. (B) Qubit spectrum and master equation simulation for a qubit drive power of -6.0 dBm. (C) Qubit spectrum and master equation simulation for a qubit drive power of -3.0 dBm. The feature at $\sim 0$ MHz is an additional HBAR mode, one free-spectral range from the HBAR mode of interest for this work. The only parameter that was varied within the simulation was the room-temperature value of the qubit drive.
@@
where $\delta$ calibrates the room-temperature power to the corresponding value of $\mathcal{P}_{\rm d}$ accounting for all losses and multiplicative factors. The value of $\delta$ was determined by matching the simulation to the power-broadened qubit spectrum at multiple qubit drive powers. A set of qubit spectra is shown in Supplementary Fig. 7 where it can be seen that the master equation simulation is in excellent agreement capturing the qubit asymmetry at low power, and the power-broadened qubit spectrum at higher drive powers. It should be noted the measured transmission signal $S_{21}$ is proportional to the qubit population $\langle\hat{\sigma}_{\rm z}\rangle$ only if the appropriate signal quadrature is measured. Here, this was determined by calculating the rotation angle that minimized the signal in the out-of-phase quadrature. To confirm, we compared our single-trace two-tone data with a direct measurement of the readout resonator frequency at several qubit drive powers. At the highest drive powers, we observed a slight deviation resulting from signal mixing into the out-of-phase quadrature. However, this was at the drive powers above the self-quenching threshold and resulted in a slight mismatch between our experimental and simulated two-tone traces.
@@
SUPPLEMENTARY FIGURE 8: Gated two-tone qubit spectroscopy. (A) Experimental gated two-tone measurements for qubit drive powers -12.0, -6.0, -3.0 dBm. The master equation simulations at the same drive power are plotted as dashed black lines. The simulation has an excellent agreement for the ringdown data for all powers and a slight deviation from the ring-up data. This could be attributed to higher-order non-linearities we are not considering within our simple master equation model. However, the agreement between the experiment and simulation is good for all powers. (B) Simulated phonon state population $\langle\hat{b}^{\dagger}\hat{b}\rangle$ as a function of qubit drive power. (C) Simulated second-order phonon correlations $g^{(2)}(0)$ as a function of qubit drive power.
@@
Following the calibration of the qubit spectrum, the phonon mode was included in the master equation simulation. The value of the coupling rate $g_{\rm qb}$ and the phonon linewidth $\gamma_{\rm b}$ were determined by performing a fit to the time domain gated two-tone measurements. The fit was determined by fitting both the ringdown (as shown in the main text) and also the ring-up of the qubit. Here, the device was allowed to thermalize before the gated-ringdown measurement, weakly probing the readout resonator continually. At the time $t=0$, the qubit drive is switched on for $175\mu$s, allowing the phonon moded to ring-up. At $t=175\mu$s, the qubit drive is switched off, and the phonon ringdown is observed, as described in the main text. The best-fit values were compared at multiple drive powers; see Supplementary Fig. 8(a). We extracted a value of the qubit phonon coupling of $g_{\rm qb}/2\pi=162$ kHz and a phonon linewidth of $\gamma_{\rm b}/2\pi=6.81$ kHz. From this set of simulations, the steady-state phonon population and phonon statistics can be estimated for multiple drive powers. We observe a good agreement between the numerical simulation and the ring-up data and an excellent agreement between the ringdown data for all powers. Deviations in the ring-up simulations likely result from higher-order nonlinearities we are not considering in our model.
@@
The master equation simulations directly calculate the time dynamics of $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$, this value has been scaled by the same constant factor for all powers to give a direct comparison to our two-tone measurement. Since we expect the variation in the transmission coefficient measured in the two-tone measurement to be proportional to $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$, see Eq. S3 and Ref. Gambetta et al. (2006). From our simulation, we calculate the phonon population and second-order correlation function for multiple drive powers, shown in Supplementary Fig. 8(b,c). For low drive powers, the phonon statistics are described by a coherent state, $g^{(2)}(0)\approx 1.0$. However, as discussed in the main text, the phonon amplitude exhibits self-quenching above a given upper-threshold power. This can be seen in both the phonon population, as a rapid decrease in the populations, and in the phonon statistics as $g^{(2)}(0)>1.0$ for drive powers above the self-quenching threshold.
SUPPLEMENTARY TABLE 1: Symbols and parameters
@@
III Supplementary Note 3: Simplified Model